OpenStudy (anonymous):

Solve for x: 1/2(4 + x) + 2 = 16

OpenStudy (anonymous):

x=24

OpenStudy (anonymous):

\[\frac{1}{2}(4+x)+2=16\] double both sides to get rid of annoying fraction \[4+x+4=32\] \[x+8=32\] \[x=32-8=24\]

OpenStudy (anonymous):

voteif its right